OpenStudy (anonymous):

Use scientific notation to divide the following two numbers. Express the answer using scientific notation; ret? (6.8*10^6)/(3.99*10^4)

OpenStudy (anonymous):

(6.8*10^6)/(3.99*10^4) = 6.8/3.99 * 10^2 = 1.704 * 10^2 Answer: 1.704*10^2

OpenStudy (anonymous):

1.7*10^2

OpenStudy (anonymous):

so it should be 1.7

OpenStudy (anonymous):

--Divide the numbers before the 10 x # by each other --Subtract the exponents of the 10s Answer should be 1.7e2 or 1.7*10^2

OpenStudy (anonymous):

it should be 1.7*10^2

OpenStudy (anonymous):

so not 1.704

OpenStudy (anonymous):

well, it depends on what you teacher or the book wants. Do you have answers to choose from? or are you told where to round up too?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

then i would go with 1.704*10^2